punisher432:

The formula for the circumference of a circle is 4072-01-01-03-00_files/i0150000.jpg. Determine the circumference when the radius r is 10 cm

Eiwoh2:

@dude @SmokeyBrown

Eiwoh2:

Also, could you save the picture or formula in the question and attach the file? We can't really see what "4072-01-01-03-00_files/i0150000.jpg" is.

punisher432:

The formula for the circumference of a circle is c=2r. determine the circumference when the radius is 10 cm

dude:

Substitute the radius into the equation \(C=2\pi 10\) Solve

punisher432:

20?

punisher432:

would it be 20 cm or 20 that symbol you used cm

dude:

\(C=2\pi 10\) =\(C=20\pi\) \(20\pi \approx 62.83\)

punisher432:

so it is 20 cm not 20 that symble cm?

dude:

is it asking in terms of pi?

punisher432:

a: 100 that symbol u used cm b: 20 same symbol cm c:20 cm d: 100 cm

dude:

20 pi \(20\pi\)

punisher432:

ok thank you so much

punisher432:

What will the temperature be in degrees Fahrenheit when it is 80° Celsius outside? (Recall the formula F = (9C 4072-01-01-03-00_files/i0190000.jpg5) + 32).

dude:

what is the image also make a new post

punisher432:

ooops ok thnks dude you rock